### v2

# LDI: Localized Data Imputation for Text-Rich Tables

Abstract.
Missing values are pervasive in real-world tabular data and can significantly impair downstream analysis. Imputing them is especially challenging in text-rich tables, where dependencies are implicit, complex, and dispersed across long textual fields. Recent work has explored using Large Language Models (LLMs) for data imputation, yet existing approaches typically process entire tables or loosely related contexts, which can compromise accuracy, scalability, and explainability. We introduce LDI, a novel framework that leverages LLMs through localized reasoning—selecting a compact, contextually relevant subset of attributes and tuples for each missing value. This targeted selection reduces noise, improves scalability, and provides transparent attribution by revealing which data influenced each prediction. Through extensive experiments on real and synthetic datasets, we demonstrate that LDI consistently outperforms state-of-the-art imputation methods, achieving up to 8% higher accuracy with hosted LLMs and even greater gains with local models. The improved interpretability and robustness also make LDI well-suited for high-stakes data management applications.
Data Imputation, Dependency Evaluation, Explainability
1. Introduction
Missing values are a common characteristic of many real-world datasets, particularly those derived from open or heterogeneous sources. If left unaddressed, they can significantly undermine the reliability of downstream analysis and machine learning applications. For example, many learning algorithms operate under the assumption that all input data is fully observed (Emmanuel et al., 2021; Zhang et al., 2008; Kumar et al., 2017), and their performance often degrades when faced with incomplete records. This not only reduces predictive accuracy but can also introduce systematic bias into analytical outcomes (Kang, 2013; Zhang et al., 2007; Qin et al., 2007; Stoyanovich et al., 2020). As a result, predicting or imputing missing values is a critical component of the broader data wrangling process, often helping to preserve the underlying structure of the data and ensuring the validity of subsequent analysis (Song and Sun, 2020).
Motivation. While the problem of data imputation has been studied for decades, a growing fraction of modern datasets are text-rich tables that mix structured fields with lengthy textual descriptions, reviews, or metadata. These tables are common in domains such as e-commerce, real estate, scientific data management, and open government data. Unlike purely numerical datasets, text-rich tables contain dependencies that are implicit, complex, and often buried within long, noisy text. Identifying which textual cues are informative for imputation remains a major open challenge.
Example. Consider a restaurant table with missing values in the City column (see Figure 1). Several other attributes—including Description, Phone, and Reviews—might help infer the correct city, but their relevance varies. The Description column may contain location-related hints (e.g., “Modern American dishes” and “Elegant Italian dining”) that are typically too general or vague to be useful, while the Reviews field may offer more specific but noisy references (e.g., “Best gelato in Vegas”). The Phone column, in contrast, encodes reliable geographic signals via area codes, though inconsistencies in formatting can obscure them. Determining which attributes and records truly contribute to accurate imputation requires both contextual understanding and selective reasoning.
Figure 1. Example of data imputation
Limitations of Existing Work. Existing approaches to data imputation range from simple heuristics to advanced generative models. Heuristic techniques (e.g., mean or mode imputation) are fast but ignore inter-attribute dependencies, often leading to biased estimates (Baraldi and Enders, 2010; Khan and Hoque, 2020). Model-based methods, including regression, expectation-maximization (EM), and matrix factorization, capture statistical dependencies but can be computationally expensive and rely on strong assumptions about the data distribution (Little and Rubin, 2019). Recent deep learning–based approaches, such as autoencoders, GANs, and transformer models, have shown promise in capturing nonlinear patterns, especially in large or multi-modal datasets (Yoon et al., 2018). Building on this progress, large language models (LLMs) have emerged as powerful tools for data imputation, thanks to their ability to model complex dependencies and generalize across diverse contexts (Narayan et al., 2022; Zhang et al., 2024; Li and Döhmen, 2024). However, current LLM-based approaches remain fundamentally global in scope as they tend to process entire tables or large portions thereof. This introduces three major issues that remain largely overlooked: (1) Scalability: Tables often exceed model context windows, forcing either truncation or splitting that loses important structure. (2) Noise sensitivity: Including loosely related rows and columns dilutes the signal, making imputations less accurate. (3) Lack of transparency: When the entire table is fed to an LLM, it becomes unclear which attributes or tuples influenced each prediction, undermining interpretability and user trust.
These limitations raise an important but underexplored question: Can missing values be accurately imputed using only a localized subset of the table—one that preserves essential context while improving scalability and explainability?
Our Approach. We hypothesize that, in many cases, missing values can be imputed accurately using a localized context—a compact subset of attributes and tuples most relevant to the missing value. Our approach, LDI, operationalizes this idea by decomposing the problem into two sub-tasks: (1) selecting a subset of columns that are most relevant to the column with missing values, and (2) identifying a subset of rows that provide sufficient contextual evidence for imputation. This localized reasoning enables LDI to reduce noise, scale to larger tables, and trace which data influenced each prediction.
Our extensive evaluation across multiple datasets demonstrates that LDI not only improves performance but also enhances interpretability and robustness, particularly in text-rich, noisy, and heterogeneous environments.
Contributions: Our contributions can be summarized as follows: (1) We introduce LDI, a framework that performs attribute- and tuple-level selection to guide LLM-based imputation. (2) By limiting context to relevant subsets, LDI achieves both higher accuracy and transparent attribution. (3) LDI handles textual inconsistencies and heterogeneous formats more effectively than existing models. (4) The framework supports small, local LLMs, enabling efficient and privacy-preserving imputation. (5) Through analytical and experimental evaluations on real-world datasets from various domains, we demonstrate that LDI achieves superior accuracy, scalability, and explainability compared to state-of-the-art methods. (6) To support reproducibility and foster further research, we publicly release our code, framework, and datasets [^1] https://github.com/soroushomidvar/LDI.
2. Related Works
Our work is closely related to two main areas of research: (1) data imputation and (2) dependency detection.
2.1. Data Imputation
Missing data imputation is a well-researched field, with a wide range of methods developed to address the challenges of handling incomplete datasets.
Traditional Methods: Basic approaches replace missing values with fixed constants, such as zero or the mean, or use strategies like Last Observation Carried Forward (LOCF) and Next Observation Carried Backward (NOCB) (Engels and Diehr, 2003; Lachin, 2016; Klamroth-Marganska et al., 2014). These approaches are computationally inexpensive and straightforward to implement but rely on restrictive assumptions, such as the randomness of missing values or ordering of tuples, that often lead to biased results (Molnar et al., 2008), limiting their applicability in real-world scenarios where data is complex and irregular.
Statistical Models: While basic methods rely on direct substitution, statistical models enhance imputation accuracy by estimating the underlying data distribution and predicting missing values accordingly. Examples include approaches based on relational dependency networks (Mayfield et al., 2010), factor graphs (Rekatsinas et al., 2017; Wu et al., 2020), distance likelihood maximization (Song and Sun, 2020), low-rank matrix completion (Candes and Recht, 2012), and non-linear models such as tree-based ensembles (Stekhoven and Bühlmann, 2012; Chen and Guestrin, 2016). While effective in certain scenarios, these approaches often rely on assumptions about data properties, such as normality or linearity, which can limit their applicability to real-world datasets.
Similarity-based Approaches: Another line of research focuses on similarity-based techniques, where missing values are inferred based on similar neighbors with complete values. Neighbor-based methods, such as k-Nearest Neighbors (KNN) (Altman, 1992) and its extension, KNN Ensemble (KNNE) (Domeniconi and Yan, 2004), find the nearest neighbors with the smallest distance on complete attributes and aggregate their values as the imputation for the missing value. Similarly, MIBOS (Wu et al., 2012) imputes missing values through a similarity-based iterative approach that counts matching attribute values. Instead of relying on individual neighbors, some methods group similar tuples into clusters and perform the imputations within each cluster. For example, CMI (Zhang et al., 2008) imputes missing values by first clustering the dataset and then generating plausible values using similar tuples in the cluster. MGI (Yan et al., 2015) applies the Gaussian mixture model (GMM), and IFC (Nikfalazar et al., 2017) uses iterative fuzzy clustering to integrate mean imputation, hot deck, and multiple imputation. However, these similarity-based methods are constrained by their similarity functions, cannot handle semantic similarity, and have limited performance if used directly on the data without additional steps.
Generative Models: Beyond these traditional methods, recent advances in generative models have opened new possibilities for data imputation by enhancing the ability to learn data patterns and generate more realistic values. For instance, GAIN (Yoon et al., 2018) uses the Generative Adversarial Nets (GAN) framework, where a generator imputes missing components and a discriminator ensures the generated data follows the true distribution. MIDA (Gondara and Wang, 2018) proposes a multiple imputation model using deep denoising autoencoders, and HI-VAE (Nazabal et al., 2020) adapts variational autoencoders to work with heterogeneous data. IPM (Mei et al., 2021) leverages pre-trained language models to capture semantics for predicting missing values in databases. It introduces IPM-Multi, which models imputation as multiclass classification, and IPM-Binary, which uses binary classification for candidate selection. LakeFill (Yang et al., 2025) extends the role of LLMs in imputation by incorporating external data lakes and focusing on tuple-level retrieval, addressing cases where local redundancy is insufficient. ZeroEC (Wu et al., 2025) focused on error correction using structured fields through LLM-based rule generation. UnIMP (Wang et al., 2025) introduces a framework for imputing mixed-type data by integrating LLMs with a message-passing network, and FMW (Narayan et al., 2022) also demonstrates that LLMs can handle missing data imputation effectively by framing them as prompting tasks. Although these models achieve state-of-the-art results, their black-box nature lacks explainability, making it difficult to understand how and why specific imputations are generated. The inability to trace an imputed value back to a specific part of the input data prevents validating, correcting, and refining imputations and limits the model’s applicability where understanding why a particular value was imputed can be just as important as the imputed value itself.
2.2. Dependency Detection
In the context of attribute selection, our method is closely related to research on identifying dependencies between attributes, particularly through the use of Functional Dependencies (FDs) and Approximate Functional Dependencies (AFDs). FDs describe exact relationships where the value of one attribute is fully determined by another. Formally, an FD $X\rightarrow A$ holds in a relation $r$ if for all tuples $t$ and $u$ in $r$, $t[X]=u[X]$ implies $t[A]=u[A]$. FDs are widely used in database design to capture exact deterministic relationships (Papenbrock and Naumann, 2016; Bohannon et al., 2006; Salem and Abdo, 2016; Al-Janabi and Janicki, 2022; Ni et al., 2024). In practice, however, FDs are often too strict due to inconsistencies or noise in real-world data. To address this, AFDs have been introduced (Kivinen and Mannila, 1995; Mandros et al., 2017). An AFD $X\rightarrow A$ holds approximately if it is satisfied by most of the data, with the quality measured by the error $e(X\rightarrow A)$, defined as the minimum fraction of tuples that must be removed to make the dependency exact (King and Legendre, 2003).
FDs and AFDs underpin many data cleaning techniques, including sparse regression for dependency discovery (Guo and Rekatsinas, 2019), scalable FD-based repairs for iterative data pipelines (Rezig et al., 2021), conditional FDs (Bohannon et al., 2006), and dynamic AFDs that tolerate dirty data by incorporating data source information (Ding et al., 2024). RENUVER (Breve et al., 2022) and Derand (Song et al., 2020) similarly leverage FDs to guide candidate selection for imputation. While effective on structured fields, these methods struggle with patterns embedded in free-text attributes and rely mainly on surface-level similarity rather semantic reasoning. They also assume consistent formatting and uniform representation, limiting their ability to capture dependencies expressed through partial matches or semantically related phrases. LLMs, in contrast, can model such dependencies more flexibly by drawing on their contextual understanding of language.
3. Preliminary
We formalize the problem of missing data imputation and outline the setting for LDI.
Problem Statement. Let $\mathcal{D}$ be a dataset with schema $\mathcal{R}$ consisting of attributes $A_{j}\in\mathcal{R}$. Each tuple $t_{i}\in\mathcal{D}$ contains a set of cell values, where a missing entry in the $j$-th attribute is denoted by $t_{i}[A_{j}]=\varnothing$. Let $\hat{\mathcal{D}}$ denote the imputed version of $\mathcal{D}$ such that, for all $i,j$ where $t_{i}[A_{j}]$ is missing in $\mathcal{D}$, the corresponding imputed tuple $\hat{t_{i}}\in\hat{\mathcal{D}}$ satisfies $\hat{t_{i}}[A_{j}]\neq\varnothing$. We denote by $\mathcal{D}^{*}$ the ground truth dataset in which each tuple $t_{i}^{*}\in\mathcal{D}^{*}$ contains the true values $t_{i}^{*}[A_{j}]$ for all attributes $A_{j}\in\mathcal{R}$.
Goal. Given a dataset $\mathcal{D}$, the objective is to produce an imputed dataset $\hat{\mathcal{D}}$ by predicting missing entries of a target attribute $A_{T}$, reconstructing the ground truth dataset $\mathcal{D}^{*}$. LDI focuses on identifying the most relevant tuples and attributes for each imputation task and leveraging the reasoning capabilities of LLMs to achieve the following: (1) high imputation accuracy in real-world datasets, even when missing values depend on complex and unknown relationships; (2) improved interpretability by tracing which parts of the input influenced each prediction; (3) robustness to noise and formatting inconsistencies without relying on strict assumptions; and (4) minimal dependence on external data or task-specific resources, enabling broader and more efficient applicability.
Assumptions. We assume that (1) some values of $A_{T}$ (the attribute with missing entries) appear more than once across tuples, allowing the model to learn associations from these repetitions, and (2) one or more textual or categorical attributes are correlated with $A_{T}$ through textual, structural, or contextual cues. Similar assumptions are commonly made in prior work on data imputation (Rekatsinas et al., 2017; Breve et al., 2022; Rezig et al., 2021; Zhang et al., 2008).
4. Proposed Method
While LLMs’ strong performance on a range of table reasoning tasks makes them promising candidates for data imputation, they often struggle with scalability to large tables (Chen, 2023; Ji et al., 2023) and lack interpretability in their imputation decisions. To address these issues, LDI, selectively identifies the most relevant attributes and tuples, focusing the model’s attention and improving explainability. In the first phase ($\S$4.1), LDI retains only the attributes that are likely related to the target attribute $A_{T}$ and informative for imputation. The second phase ($\S$4.2) filters tuples to produce a diverse and relevant set of records. Finally, in the third phase ($\S$4.3), the selected attributes and tuples are fed into the LLM, which performs the imputation. An overview of this process is demonstrated in Figure 2.
Figure 2. Overview of the LDI framework
4.1. Attribute Selection
To identify attributes most relevant to the imputation target, LDI performs attribute selection in two main phases: (1) detecting group-specific patterns in candidate columns, and (2) evaluating approximate dependency relationships.
While traditional functional dependencies can model relationships between attributes, they are often too strict for real-world tables that contain noise and long textual values. To better handle such cases, we introduce a relaxed dependency criterion tailored to noisy, text-rich data.
Definition 0 ((p, q)-Approximate Dependency).
Let $X$ and $A$ be columns in a table, and let $p,q\in[0,1]$. We say that a $(p,q)$-approximate dependency $X\rightarrow A$ holds if, for at least a fraction $p$ of the distinct values $\alpha$ in column $A$, there exists a common substring $s$ in column $X$ such that:
  • The probability that $s$ appears in a record given the value $\alpha\in A$ is at least $q$, i.e., $P(s\in X\mid A=\alpha)\geq q$, and
  • For all $\alpha^{\prime}\neq\alpha$, $P(s\in X\mid A=\alpha^{\prime})<q$.
This relaxed definition captures partial and group-specific relationships between attributes. It relies on substring-level evidence, making it robust to formatting inconsistencies and local variations. The parameters $p$ and $q$ allow the strictness of the dependency to be tuned, enabling a balance between robustness and sensitivity to weak signals.
We use this criterion to identify a subset of candidate attributes $A_{C}$ that exhibit a $(p,q)$-approximate dependency with the target attribute $A_{T}$ with the caveat that such dependencies, like traditional functional dependencies, cannot be definitively inferred from a database instance, and one can only verify that a possible dependency is not violated.
4.1.1. Pattern-Based Dependency Detection
To assess whether a candidate attribute $A_{C}$ exhibits a dependency on the target attribute $A_{T}$, LDI analyzes its values for group-specific patterns. We use the Longest Common Substring (LCS) to detect recurring character sequences—such as prefixes, category tags, or embedded keywords—that frequently appear within groups sharing the same $A_{T}$ value. These patterns provide localized signals that are informative for imputation, even if they do not hold globally.
Within each group (i.e., tuples sharing the same $A_{T}$ value), we extract all substrings from $A_{C}$. A substring is retained if it appears in at least a $q$ fraction of the group’s tuples, therby filtering out infrequent or incidental patterns while balancing robustness to noise with sensitivity to meaningful regularities. Each frequent substring in $A_{C}$ must be uniquely associated with a single value in the $A_{T}$ column; otherwise, it violates the functional dependency condition. To enforce this constraint, we discard any substring that appears as frequent in more than one group.
Additionally, we remove redundant substrings that are fully contained in longer, more informative ones within each group. Since every frequent substring implies that all of its subparts are also frequent, retaining them only increases clutter. While this cleanup step is optional and does not affect the correctness, it simplifies interpretation and highlights the most distinctive patterns for explainability.
For example, consider the Phone attribute in our running example (Figure 1). Las Vegas phone numbers (e.g., 702/876-4190, 702/731-4036, +1702-385-7111) share a common prefix pattern, whereas New York numbers (212/333-7788, 212/431-0021, 212/688-5447) follow a different pattern. With a threshold $q=0.6$, substrings such as $\{\mbox{‘702/’},\mbox{‘702’},\mbox{‘02/’},\mbox{‘70’},\ldots\}$ are retained for Las Vegas, and $\{\mbox{‘212/’},\mbox{‘212’},\mbox{‘12/’},\mbox{‘21’},\ldots\}$ for New York. Substrings that appear across groups or are fully contained within longer patterns are discarded, leaving only the most distinctive patterns (e.g., ‘702/’ for Las Vegas and ‘212/’ for New York).
The LCS method is particularly effective in uncovering non-trivial patterns, even in the presence of noises such as typos or formatting inconsistencies. Its robustness stems from two key properties: (1) If a pattern is missing in some cells of a group, the LCS across the remaining cells is preserved. Adjusting the parameter $q$ makes it possible to relax the matching requirement while maintaining the detection of meaningful dependencies. (2) Inconsistent formatting produces shorter but still relevant LCS results. For example, variations such as ‘+1780 ’, ‘780/’, and ‘780-’, strill reveal the underlying pattern ‘780’.
4.1.2. Dependency Evaluation
In the second step of the attribute selection phase, we evaluate whether a $(p,q)$-approximate dependency exists between $A_{C}$ and $A_{T}$. Specifically, the algorithm counts how many groups have at least one unique pattern—i.e., a substring that frequently appears within that group but is uncommon across others. If such patterns are found in at least a fraction $p$ of all groups, we consider $A_{T}$ to be approximately dependent on $A_{C}$. This criterion ensures that the inferred dependency is supported by consistent evidence across multiple groups, rather than arising from isolated or random occurrences.
[…]
This process is applied to all candidate attributes, and those satisfying the $(p,q)$-dependency condition are retained for the next phase.
4.1.3. Scalability through Stratified Group Sampling
Although our method can efficiently detect dependencies on large datasets (see $\S$5.1 and $\S$6.6), datasets with hundreds of thousands or millions of records may still require additional measures to ensure scalability. To this end, we introduce an optional group sampling step.
While LDI can operate on a random sample of tuples, naive random sampling is inadequate as it overrepresents frequent values in $A_{T}$ and underrepresents rare ones, which can obscure group-level patterns. Since LDI relies on identifying such patterns, it is essential to preserve the diversity of $A_{T}$ in the sample.
Group Sampling. To maintain diversity in the target attribute $A_{T}$, we partition the table by distinct values of $A_{T}$ and then sample in a structured way. We randomly select $m$ groups that each contain at least $n$ rows, and within each group, we randomly draw $n$ rows, producing a sample of $m\times n$ tuples.
When the dataset lacks sufficient tuples for a full $m\times n$ sample, such as when some values of $A_{T}$ are infrequent, we include as many complete groups as possible, prioritizing those with larger sizes. Specifically, we first select all groups with at least $n$ rows, then add those with n1n-1 rows, continuing until no further groups can be added without exceeding the sample limit.
The parameters $m$ (number of groups) and $n$ (tuples per group) control the breadth and depth of the sample. Increasing $m$ covers more distinct $A_{T}$ values, while increasing $n$ captures finer-grained or noisier within-group patterns. Smaller values of $(m,n)$ may suffice for more homogeneous datasets.
This stratified approach preserves group-level diversity while keeping the sample size manageable, enabling LDI to scale effectively without sacrificing the reliability of dependency detection.
4.2. Tuple Selection
In the second phase of our proposed method, we aim to identify a small set of tuples to serve as examples for guiding the LLM during imputation. The selection process follows three key criteria: (1) tuples must have complete values in $A_{T}$, (2) they should be similar to the tuple with the missing value to provide relevant context, and (3) they should be diverse in their $A_{T}$ values to reduce the risk of bias.
We first filter the dataset to retain only tuples with complete values in $A_{T}$, as they are needed to serve as valid examples [^2] Tuples may have missing values in other attributes since full completeness is not required, but they are less likely to be similar to the selected tuple for imputation.. To avoid distractions from irrelevant attributes, similarity is computed over the subset of relevant attributes identified in the previous phase. For a tuple $t_{i}$ with a missing $A_{T}$ value, its similarity to a complete tuple $t_{j}$ is measured using the normalized length of their LCS across the selected attributes $S$:
(1) $\text{similarity}(t_{i},t_{j})=\frac{1}{|S|}\sum_{A\in S}\frac{|\text{LCS}(t_{i}[A],t_{j}[A])|}{\max(|t_{i}[A]|,|t_{j}[A]|)}$
Finally, we select the top $k$ tuples with distinct values in $A_{T}$, balancing similarity with diversity to ensure the examples are both relevant and representative. These $k$ examples are then passed to the LLM for few-shot imputation.
Suppose the relevant attribute selected in the first phase for our running example is Phone, and we want to impute the missing City for tuple $t_{i}$ corresponding to Moongate in Figure 1. The following table presents the selected subset:
Tuple Phone City
$t_{1}$ 212/333-7788 New York
$t_{2}$ 702/876-4190 Las Vegas
$t_{3}$ 212/431-0021 New York
$t_{4}$ 702/731-4036 Las Vegas
$t_{5}$ +1702-385-7111 Las Vegas
$t_{6}$ 212/688-5447 New York
$t_{i}$ 702-791-7352 N/A
We compute the pairwise similarity of $t_{i}$ with each complete tuple based on their Phone values as follows:
Pair LCS Length Max Length Similarity
$t_{i}$, $t_{1}$ -7 2 12 0.167
$t_{i}$, $t_{2}$ 702 3 12 0.250
$t_{i}$, $t_{3}$ 02 2 12 0.167
$t_{i}$, $t_{4}$ 702 3 12 0.250
$t_{i}$, $t_{5}$ 702- 4 14 0.286
$t_{i}$, $t_{6}$ 7 1 12 0.083
Ranking the tuples by similarity gives:
  1. (1)
    $t_{5}$: 0.286 (City = Las Vegas)
  2. (2)
    $t_{2}$: 0.250 (City = Las Vegas)
  3. (3)
    $t_{4}$: 0.250 (City = Las Vegas)
  4. (4)
    $t_{1}$: 0.167 (City = New York)
  5. (5)
    $t_{3}$: 0.167 (City = New York)
  6. (6)
    $t_{6}$: 0.083 (City = New York)
When $k=1$, the most similar tuple $t_{5}$ is selected. When $k=2$, the algorithm chooses $t_{5}$ and $t_{1}$ to preserve similarity while covering different City values.
4.3. Data Imputation Using LLM
In the third phase of LDI, we leverage LLMs to impute missing values using few-shot learning. The model is provided with a small set of example tuples that include the relevant attributes identified in § 4.1 and retrieved similar tuples from § 4.2. The LLM then infers the missing value by recognizing patterns across these examples.
Following prior work (Narayan et al., 2022), each tuple is serialized as a sequence of key-value pairs. The input prompt consists of a task description, example tuples, and a query tuple with a missing value:
[Context]
  A short description of the task.
[Examples]
  Example 1:
  Attr. A: Val. A1, Attr. B: Val. B1, ...
  Target Attr.: Target Val. 1,
  Example 2:
  Attr. A: Val. A2, Attr. B: Val. B2, ...
  Target Attr.: Target Val. 2,
[Query]
  Attr. A: Val. Ax, Attr. B: Val. Bx, ...
  Target Attr.: ?
The LLM leverages the provided examples to predict the missing value for the target attribute $A_{T}$, enabling it to capture complex dependencies, handle noisy data, and infer plausible values even for patterns not seen in the examples.
LLMs offer several key advantages in this imputation task. First, they demonstrate strong semantic understanding by leveraging extensive pretraining to infer values based on real-world knowledge. For example, they can associate area codes with their corresponding cities or map currency codes to countries, all without requiring external lookup tables. Second, LLMs are capable of generalizing to unseen cases: even when a specific pattern has not appeared in the provided examples, the model can infer plausible values by drawing on its understanding of similar contexts. Finally, LLMs exhibit robustness to noise and formatting inconsistencies. They can recognize equivalences among different representations of the same entity—for instance, interpreting ‘+1-702’, ‘702/’, and ‘702-’ as referring to the same area code—allowing them to handle noisy or variably formatted data with ease.
Running Example Recap. Suppose we aim to impute the missing City value for the tuple $t_{i}$ corresponding to Moongate in Figure 1. The method first groups the data by City and samples $m$ groups with $n$ tuples each (Figure 3). Pattern detection identifies recurring substrings in candidate columns. Here, both Phone and Reviews show repeated substrings, but only Phone satisfies the group-level dependency criteria: area codes are unique to cities, whereas patterns in Reviews (e.g., “[(’Rated ”) appear across multiple cities and are discarded. Next, similarity search identifies $k$ tuples most relevant to $t_{i}$ based on the selected attribute (Phone) while ensuring diversity in City values. These tuples are serialized into key-value format and provided to the LLM in a few-shot setup. The model then imputes the missing city by relying solely on informative signals in the Phone column, producing predictions that are both accurate and traceable. This workflow demonstrates how LDI reduces noise, focuses on informative data, and leverages LLM reasoning to produce explainable imputations, even without predefined dependencies in the dataset.
Figure 3. Example of data imputation (after applying our approach)
5. Performance Analysis
In this section, we analyze the computational cost of the main components of our method. We first present supporting lemmas for key steps in each phase.
Lemma 0.
Given a collection of $n$ strings, each of length $\ell$, all substrings that appear in at least $q$ fraction of the strings can be found in time $\mathcal{O}(n\cdot\ell+r)$ and space $\mathcal{O}(n\cdot\ell)$, where $r$ is the number of substrings reported.
Proof.
We build a Generalized Suffix Tree (GST) for the $n$ input strings, each of length $\ell$, after appending a unique terminal symbol to each string. This tree can be constructed in $\mathcal{O}(n\cdot\ell)$ time and space using Ukkonen’s algorithm (Ukkonen, 1995). Each leaf is labeled with its corresponding string ID, and a post-order traversal is used to annotate each internal node with the set of string IDs in its subtree. For each node, if the number of distinct string IDs is at least $\lceil qn\rceil$, we report the substring represented by that node. Reporting all such substrings takes $\mathcal{O}(r)$ time, where $r$ is the number of substrings found. Therefore, the total time and space complexity is $\mathcal{O}(n\cdot\ell+r)$ and $\mathcal{O}(n\cdot\ell)$, respectively. ∎
Lemma 0.
Given a collection of $s$ strings, each of length $\ell$, partitioned into $m$ groups, the time to determine whether each string appears uniquely within a single group is $\mathcal{O}(s\cdot\ell)$.
Proof.
To determine whether each string appears in only one group, we use a hashmap that associates each string with the set of group identifiers in which it appears. We iterate over all $s$ strings and, for each one, insert it into the hashmap and record its group identifier. Since each string has length $\ell$, inserting or comparing a string in the hashmap takes $\mathcal{O}(\ell)$ time. Processing all $s$ strings therefore takes $\mathcal{O}(s\cdot\ell)$ time. After building the hashmap, we scan its entries to check whether any string is associated with more than one group. As there are at most $s$ unique strings, this step takes $\mathcal{O}(s)$ time. Hence, the overall time complexity is $\mathcal{O}(s\cdot\ell)$. ∎
5.1. Attribute Selection
The time complexity of the attribute selection phase depends on the number of candidate attributes ($a_{c}$), the number of groups formed by values in the target attribute ($m$), the number of sampled tuples in each group ($n$), and the average string length in attribute values ($\ell$). The algorithm consists of four steps: (1) an optional group sampling, (2) pattern detection, (3) uniqueness checking across groups, and (4) an optional redundancy filtering step.
In the first step, the algorithm samples a subset of the dataset by grouping tuples according to their values in the target attribute $A_{T}$. It then randomly selects $m$ groups and samples $n$ tuples from each group. Letting $|\mathcal{D}|$ be the size of the original dataset, this grouping and sampling process takes $\mathcal{O}(|\mathcal{D}|)$ time, as each tuple is processed once.
In the second step, the algorithm detects common substrings that appear in at least $q$ fraction of the tuples within each group. This is done using a GST, which enables efficient substring frequency analysis. According to Lemma 1, the GST can be built in $\mathcal{O}(n\cdot\ell)$ time per group, and all frequent substrings can be extracted in total time $\mathcal{O}(n\cdot\ell+r)$, where $r$ is the number of substrings reported. Applying this across $m$ groups and $a_{c}$ attributes results in a total time complexity of $\mathcal{O}(a_{c}\cdot m\cdot(n\cdot\ell+r))$.
In the third step, the algorithm checks whether each substring appears uniquely within a single group. Lemma 2 shows that this can be efficiently done using a hashmap that associates each substring with the set of groups it appears in. Letting $s=m\cdot r$ be the total number of substrings across all groups, this step takes $\mathcal{O}(s\cdot\ell)$ time per attribute. Across all $a_{c}$ attributes, the total cost becomes: $\mathcal{O}(a_{c}\cdot m\cdot r\cdot\ell)$.
An optional fourth step removes redundant substrings that are fully contained within longer ones in each group. This improves the clarity and specificity of the selected patterns but does not affect the correctness of the algorithm. This step can be done using pairwise comparisons among the $r$ substrings in each group. The cost of this step is $\mathcal{O}(r^{2}\cdot\ell)$ per group. Across all $m$ groups and $a_{c}$ attributes, this adds an additional cost of $\mathcal{O}(a_{c}\cdot m\cdot r^{2}\cdot\ell)$.
Combining all steps, the total time complexity of the attribute selection phase depends on whether the optional redundancy removal step is applied.
Without the optional step, the total time complexity of the attribute selection phase is:
$\mathcal{O}(|\mathcal{D}|+a_{c}\cdot m\cdot\ell\cdot(n+r))$
With the optional redundancy removal step, the total time complexity becomes:
$\mathcal{O}(|\mathcal{D}|+a_{c}\cdot m\cdot\ell\cdot(n+r^{2}))$
when the optional step is enabled, we incur an additional cost of $r^{2}\cdot\ell$ per group to remove substrings contained in longer ones.
5.2. Tuple Selection
To impute a missing value in a tuple $t_{i}$, our method selects $k$ similar and diverse examples from the dataset. Let $c$ be the number of complete tuples (i.e., tuples with known values in $A_{T}$), and let $a_{s}$ be the number of selected attributes. For each attribute $A$, we compute the similarity between $t_{i}[A]$ and $t_{j}[A]$ for all $c$ complete tuples $t_{j}$.
To do this efficiently, we build a GST over the $c+1$ strings (one from $t_{i}[A]$ and $c$ from complete tuples) for each selected attribute. Building the GST takes $\mathcal{O}(c\cdot\ell)$ time per attribute (see Lemma 1). After the tree is built, the LCS between $t_{i}[A]$ and any $t_{j}[A]$ can be found in $\mathcal{O}(\ell)$ time. Therefore, computing the similarity between $t_{i}$ and all $c$ complete tuples for all $a_{s}$ attributes needs a total time of $\mathcal{O}(a_{s}\cdot c\cdot\ell)$ for the similarity computation.
After calculating all similarity scores, we sort the $c$ tuples in $\mathcal{O}(c\log c)$ time to identify the top $k$ diverse most similar examples. A final scan to ensure diversity in $A_{T}$ values among the top $k$ adds $\mathcal{O}(c)$ time, which is dominated by the sorting. Hence, the total time complexity of the tuple selection phase for one incomplete tuple is:
$\mathcal{O}(a_{s}\cdot c\cdot\ell+c\log c)$
5.3. Imputation Using LLM
LDI makes one LLM call per missing value, resulting in a total number of calls equal to the number of missing entries. Each prompt includes a short fixed-length context, $k$ examples, and a query (see Figure 2). Both the query and examples in LDI are restricted to $a_{s}$ selected attributes. Assuming the context contains $\lambda$ tokens and each attribute value has an average of $\tau$ tokens, the total token usage per prompt is approximately:
$\lambda+k\times a_{s}\times\tau.$
6. Experiments
This section presents a comprehensive evaluation of our proposed data imputation framework, LDI. The experiments aim to answer three main questions: (1) How does attribute and tuple selection affect imputation accuracy and explainability? (2) How does LDI compare to existing state-of-the-art baselines? (3) How robust and scalable is LDI across datasets and missingness conditions?
6.1. Datasets
We evaluate LDI on four real-world datasets from three distinct domains, summarized in Table 1. The datasets vary widely in size, number of attributes, and the characteristics of incomplete data, providing a diverse and realistic testbed. The incomplete attributes are categorical, with vocabulary sizes ranging from 49 in the Restaurant dataset to 385 in the Phone dataset. Buy and Restaurant contain fewer attributes compared to our other datasets but are widely used in data imputation studies(Narayan et al., 2022; Chen et al., 2024; Song and Sun, 2020; Mei et al., 2021), making them valuable for comparison purposes. In contrast, Zomato and Phone datasets include a larger number of attributes, enabling a more comprehensive evaluation of attribute selection and scalability.
Table 1. Dataset summary
Dataset Domain #Rows #Attr. Incomplete Attr. (Vocabulary Size)
Buy Digital Retail 651 4 manufacturer (62)
Restaurant Food Services 864 5 city (49)
Zomato Food Services 51717 17 location (93)
Phone Mobile Devices 413840 6 brand name (385)
6.2. Experimental Setup
All experiments were conducted on an Ubuntu 20.04 machine with an AMD Ryzen 9 3900X 12-Core Processor, and 64 GB of memory, using Python version 3.9.20. For the Buy and Restaurant datasets, all available rows were utilized, while for larger datasets (Zomato, Phone), 1000 rows were randomly selected to manage the computational cost and latency of LLM calls. The same samples were reused across all methods to ensure fairness, and each experiment was repeated five times, with the averaged results reported. Group sampling was configured with $m=10$ groups and $n=10$ samples per group. In cases where the dataset had fewer groups or samples (<10<10), the maximum available was used. Thresholds $p$ and $q$ were both set to 0.9 for Buy and Phone, and 0.8 for Zomato and Restaurant, allowing slight flexibility for noisy or inconsistent datasets. A 0.9 threshold permits up to one unmatched row per group and one group with no pattern, maintaining strong dependency evidence while tolerating minor noise. For LLM-based imputation, we primarily employed GPT-4o-mini (version 2024-07-18), and also included Llama 3.2 3B to evaluate performance in limited-resource settings. While GPT-4o-mini is a larger, high-performing model, Llama 3.2 3B demonstrates LDI remains effective with smaller, open-weight models—highlighting its portability and generalizability.
6.3. Data Reduction via Attribute Selection
Table 2 highlights the efficiency of our model in reducing data size by focusing only on relevant attributes. Unlike other methods that use all available attributes, such as FMW, our approach selects a smaller set of dependent attributes identified in the initial phase. This targeted selection leads to a significant reduction in the amount of data processed, as shown by the lower character counts across all datasets. The reduction in characters is computed by comparing the average length of the selected attributes to that of all attributes. The effect is substantial across all benchmarks: in the Zomato dataset, character input is reduced by 95.7% through the selection of only 2 out of 16 attributes. Similarly, the Buy, Restaurant, and Phone achieve reductions of 57.1%, 39.9%, and 74.0%, respectively. As a result, the process becomes more explainable and the outcomes more traceable, since it is easier to understand which attributes contribute to the final results.
Table 2. Average reduction in number of attributes and characters
#Attr. (Except Target)
Dataset All Attr. (FMW) Dep. Attr. (LDI) Less Data (#Chars)
Buy 3 1 57.1%
Restaurant 4 2 39.9%
Zomato 16 2 95.7%
Phone 5 1 74.0%
6.4. Performance Compared to Baselines
We compare LDI with both traditional and LLM-based baselines: KNNE (Domeniconi and Yan, 2004) and MIBOS (Wu et al., 2012) (nearest-neighbor-based), CMI (Zhang et al., 2008) (clustering-based), ERACER (Mayfield et al., 2010), Baran (Mahdavi and Abedjan, 2020), and HoloClean (Rekatsinas et al., 2017) (statistics-based), along with two approaches that use language models: IPM (Li and Döhmen, 2024) and FMW (Narayan et al., 2022). Baran, originally an error correction model, is adapted for data imputation by replacing each missing value with a null token and predicting its correction. The number of labels (examples) is set to 20, consistent with prior studies (Li and Döhmen, 2024). To simulate missing data, we apply a missing completely at random mechanism to originally complete datasets, removing 10% of the values. Table 3 reports the accuracy of LDI compared to these baselines across four datasets. Our method achieves the highest accuracy on three datasets (Buy, Restaurant, and Phone) and the second-highest on Zomato. LDI, IPM, and FMW all utilize language models and demonstrate high performance across all datasets. This creates a clear gap between them and traditional methods, which may perform reasonably on a single dataset but fail to generalize. For instance, Baran and HoloClean perform well only on Zomato, with substantially lower results on the other datasets.
Table 3. LDI accuracy compared to baselines at 10% missing rate
Dataset KNNE MIBOS CMI ERACER Baran HoloClean IPM FMW (k=10) LDI (k=10)
Buy 0.355±0.016 0.016±0.011 0.653±0.029 0.003±0.006 0.227±0.173 0.162±0.043 0.965±0.015 0.928±0.033 0.974±0.026
Restaurant 0.206±0.019 0.093±0.025 0.560±0.049 0.185±0.031 0.315±0.044 0.331±0.056 0.772±0.023 0.766±0.122 0.832±0.054
Zomato 0.581±0.005 0.024±0.001 0.746±0.006 0.001±0.000 0.938±0.023 0.956±0.005 0.995±0.001 0.946±0.028 0.974±0.011
Phone 0.466±0.000 0.000±0.000 0.404±0.000 0.175±0.000 0.162±0.073 0.130±0.055 0.867±0.002 0.944±0.031 0.970±0.016
IPM treats imputation as a classification task and introduces two variants: IPM-Multi and IPM-Binary [^3] Under the IPM column, we report the better result of these two for each dataset.. IPM-Multi formulates the problem as multiclass classification over the candidate set, and IPM-Binary ranks each candidate using a binary classification model. Both require fine-tuning to capture the relation between context and candidate values, making the model resource-intensive and less flexible. Moreover, IPM performance is sensitive to the quality of training data and the chosen modeling strategy. In practice, large and inconsistent gaps are often observed between IPM-Multi and IPM-Binary across datasets. In contrast, LDI avoids the need for training data by dynamically selecting informative attributes and constructing few-shot prompts using relevant examples. This allows LDI to adapt to new datasets more easily and generate more explainable localized prompts that are better aligned with the LLM’s strengths. FMW performs imputation by prompting a language model with tuples from the dataset, where examples can be randomly or manually selected. All attributes are included in the prompt, regardless of their relevance, which can introduce noise and distract the model. While FMW avoids training, is simple to apply, and allows manual selection of informative examples, this selection does not scale to large datasets [^4] We therefore use random selection for comparison. and their method does not distinguish between informative and irrelevant attributes. In contrast, LDI focuses on localized context and selects only the most relevant attributes of similar tuples, leading to more accurate, traceable, and therefore explainable results across datasets.
6.5. Robustness to Different Missingness Levels
LDI is designed to maintain high accuracy regardless of the missing value rate, as long as it is provided with a few well-chosen examples. For best results, these examples should be sufficient to detect dependencies and also provide contextual information to frame the imputation task. This behavior is similar to Programming by Examples (PBE) methods, where the model generalizes well from a small, informative set of examples. To evaluate the robustness of LDI to changes in missing rate, we vary the missing rate from 10% to 50%, with a missing completely at random mechanism applied to complete tuples. As shown in Figure 4, even after imputing 50% of the data, the accuracy does not drop significantly, indicating that the model can still identify attribute dependencies and use similar examples to guide the LLM. This robustness is expected to extend to different missingness mechanisms—missing at random and missing not at random—because the model can uncover attribute dependencies even from limited samples, as reflected in the results, and a small set of informative examples is sufficient to drive accurate imputations.
Figure 4. Impact of varying missing rates on accuracy
6.6. Runtime and Scalability
Table 4 summarizes the execution time of the attribute and tuple selection phases on our real-world datasets. The attribute selection phase, executed once to identify dependent columns, completes in negligible time across all datasets. The slightly higher values in the Zomato and Phone datasets stem from their larger number of attributes and longer average string lengths. In contrast, the tuple selection phase must be performed for each missing value, and its runtime depends mainly on the number of tuples compared during similarity search, as well as the number and length of attributes selected in the first phase.
To further assess scalability with respect to input characteristics—the number of groups ($m$), tuples per group ($n$), and attribute value length in characters (ll)—we conducted controlled experiments on synthetic data. We varied $l$ from 10 to 1000 and both $m$ and $n$ from 10 to 50. As shown in Section 5, these parameters have the largest impact on the runtime of attribute selection. The number of returned substrings was controlled by setting $p=q=0.5$. The results are shown in Figure 5. As expected, runtime increases with both the number of tuples and the average string length during the attribute selection phase (Figure 5(a)), but remains practical even for lengthy strings and thousands of tuples. Group sampling can further bound this cost by operating on a small yet representative subset of the data; for instance, sampling only 100 tuples ($m=n=10$) was sufficient in our experiments to detect reliable patterns. The runtime of the tuple selection phase exhibits a similar trend (Figure 5(b)), growing moderately as the number and length of tuples increase. Since this phase must be repeated for every missing entry, its cost scales linearly with the number of imputations. When multiple attributes are selected in the first phase, the runtime increases proportionally, as similarity is computed pairwise across attributes. Nevertheless, even in these settings, runtime remains well within practical limits, supporting the scalability of our approach.
Table 4. Runtime on real-world datasets (in seconds)
Dataset Attr. Selection Tup. Selection
Buy 0.022 0.058
Restaurant 0.012 0.116
Zomato 0.338 0.375
Phone 0.795 0.162
(a) Attr. Selection
(b) Tup. Selection
Figure 5. Runtime trends on synthetic data (in seconds)
6.7. Ablation Study
To better understand the contribution of each phase of our method, we conduct a step-by-step performance evaluation by incrementally applying the attribute and tuple selection phases, starting from a method that uses all attributes and randomly selected tuples. This layered analysis allows us to observe how each phase improves the results when added on top of the other. Unlike typical ablation studies that remove one component at a time, we do not completely remove the first phase in isolation, as the second phase relies on the selected attributes from the first phase to compute similarity, and measuring it over all attributes would mask the effect of relevant signals due to noisy or redundant fields.
We evaluate performance using two metrics: exact match accuracy and ROUGE-1 F1 score. The former measures whether the model prediction is exactly correct, while the latter provides a softer measure of correctness, crediting partial matches. Typically, ROUGE is used to evaluate summarization and generation tasks, and ROUGE-1, specifically, calculates the overlap of unigrams between the prediction and the reference. In our case, using ROUGE-1 F1 helps distinguish between predictions that are completely incorrect and those that contain part of the correct value or provide a more verbose version of it. We report ROUGE-1 instead of higher-order ROUGE-N metrics, as most target values are short (1-2 tokens), making unigram overlap a more reliable measure.
6.7.1. Impact of Attribute Selection Phase
Table 5 presents results for a variant of LDI in which tuples are randomly sampled, but only attributes that are in $(p,q)$-approximate dependency with the target attribute are retained. This setting allows us to evaluate the standalone effect of removing irrelevant attributes.
Overall, by narrowing the input to the most informative data, the model can better focus on the patterns that truly matter. The results indicate that retaining only the dependent attributes preserves, and in some cases improves, the model’s performance. This shows that our attribute selection mechanism is effective at identifying the information needed for imputation, and successfully filtering out distracting or redundant features. An especially notable improvement can be observed in the Zomato dataset, which contains a larger number of attributes. For instance, in Zomato with $k=3$, exact match accuracy improves from 0.846 to 0.958 and ROUGE-1 F1 score increases from 0.866 to 0.988 when irrelevant attributes are removed. These gains are more evident when $k=3$, reflecting the importance of minimizing input noise when the number of in-context examples is limited.
In addition to the observed accuracy improvement, this phase also improves the traceability and explainability of the model. By limiting the model’s input to only the relevant attributes, it becomes clearer which evidence the model is relying on to make predictions. For instance, if the task is to impute a missing city name and the only input provided is the phone number, any correct prediction can be directly attributed to the LLM’s understanding of phone number patterns, excluding other possible cues, and if a prediction is incorrect, users can more confidently analyze whether the selected attributes were sufficient, or whether an incorrect pattern was fed. This makes the model’s decision process easier to validate and debug.
Table 5. Impact of attribute selection phase on accuracy with randomly selected tuples
(a) $k=3$
Exact Match ROUGE-1 F1
Dataset All Attr. Dep. Attr. Gain All Attr. Dep. Attr. Gain
Buy 0.923 0.923 0.0% 0.985 0.985 0.0%
Restaurant 0.747 0.761 +1.9% 0.894 0.885 -1.0%
Zomato 0.846 0.958 +13.2% 0.866 0.988 +14.1%
Phone 0.950 0.952 +0.2% 0.980 0.980 0.0%
(b) $k=10$
Exact Match ROUGE-1 F1
Dataset All Attr. Dep. Attr. Gain All Attr. Dep. Attr. Gain
Buy 0.928 0.923 -0.5% 0.985 0.985 0.0%
Restaurant 0.766 0.772 +0.8% 0.890 0.890 0.0%
Zomato 0.946 0.964 +1.9% 0.966 0.988 +2.3%
Phone 0.944 0.942 -0.2% 0.978 0.980 +0.2%
6.7.2. Impact of Tuple Selection Phase
Table 6 compares two settings: one using randomly selected tuples and the other using the proposed diverse similarity-based method. Note that in our method, the similarity between tuples is explicitly computed over the attributes identified in phase one. This ensures that the selected examples are similar in the most informative dimensions, avoiding irrelevant features that could dilute the closeness of the examples.
The results show that our similarity-based tuple selection leads to substantial improvements in exact match accuracy across all datasets and values of $k$. For example, in the Buy dataset with $k=3$, accuracy increases from 0.923 (random) to 0.974 (diverse similarity). Similar trends are observed in the Phone dataset (0.952 to 0.976) and Zomato dataset (0.958 to 0.976). Likewise, in the more difficult Restaurant dataset, the model achieves better performance, improving from 0.761 to 0.816 at $k=3$, and from 0.772 to 0.832 at $k=10$. Interestingly, while the accuracy improvements are often substantial, the changes in ROUGE scores are minimal with only minor differences across datasets. This suggests that the similarity-based selection method helps the model make more precise and fully correct predictions, as opposed to simply improving surface-level similarity. This precision stems from example relevance, where the selected examples often contain patterns or entities that reinforce the correct level of specificity. Consider the case of imputing brand names in the Buy dataset. A randomly chosen set of examples might not include any device from the LG brand. As a result, when the input tuple contains an LG product, the model may fall back to predicting the generic term “LG”, which provides only partial information. In contrast, our method is more likely to retrieve an example containing the full expression “LG Electronics”, allowing the model to generate a more specific and correct output. Similar examples can be seen with “Sony” vs. “Sony Ericsson” in Phone, “HSR” vs. “HSR Layout” in Zomato, and “New York” vs. “New York City” in Restaurant datasets.
Moreover, the tuple selection strategy implicitly leverages redundancy and repetition in the data. Many datasets contain repeated patterns, such as multiple tuples referring to the same location, brand, or product variant. By selecting a diverse yet similar set of examples, our method helps the LLM generalize more effectively from minimal context. This is particularly valuable in few-shot settings with limited in-context examples, where each example has a strong influence on the model’s output.
Table 6. Impact of the tuple selection phase on accuracy using only the dependent attributes
(a) $k=3$
Exact Match ROUGE-1 F1
Dataset Rand. Div. Sim. Gain Rand. Div. Sim. Gain
Buy 0.923 0.974 +5.5% 0.985 0.991 +0.6%
Restaurant 0.761 0.816 +7.2% 0.885 0.887 +0.2%
Zomato 0.958 0.976 +1.9% 0.988 0.983 -0.5%
Phone 0.952 0.976 +2.5% 0.980 0.986 +0.6%
(b) $k=10$
Exact Match ROUGE-1 F1
Dataset Rand. Div. Sim. Gain Rand. Div. Sim. Gain
Buy 0.923 0.974 +5.5% 0.985 0.991 +0.6%
Restaurant 0.772 0.832 +7.8% 0.890 0.889 -0.1%
Zomato 0.964 0.974 +1.0% 0.988 0.983 -0.5%
Phone 0.942 0.970 +3.0% 0.980 0.980 0.0%
The ablation results confirm that both attribute and tuple selection are critical to the effectiveness of LDI. Phase one improves the model traceability and paves the way for effective tuple selection in the next phase. Phase Two then selects diverse and similar examples that help the model better capture subtle variations in the data. Together, these phases guide the model with focused, representative context, leading to improved accuracy and more interpretable behavior.
6.8. Varying p and q
To evaluate the robustness of LDI under varying levels of noise in approximate dependencies, we conducted an experiment where p and q varied from 0 to 0.75 in increments of 0.25, and the performance was evaluated over five independent runs for each setting. The experiment was performed on the Zomato dataset, whose larger number of columns allows us to better observe how the attribute selection step responds to different noise levels. We omit results for settings where p and q exceed 0.75, as the Zomato dataset is particularly noisy and contains many inconsistencies in its fields. Applying stricter thresholds (e.g., p, q =1) would eliminate nearly all dependencies, leaving no attributes selected in many samples, even when some contain informative patterns. When either parameter is set to 0, no constraint is imposed on the dependency relationship, and all attributes are retained. In particular, setting $q=0$ removes the frequency threshold, allowing any substring that co-occurs with a distinct value in the target column to be accepted as a valid pattern. Similarly, setting $p=0$ eliminates the requirement on the minimum number of distinct target values that must support the dependency.
In Table 7, we evaluate how varying the hyperparameters $p$ and $q$ affects the imputation performance on the Zomato dataset when $k=3$ examples are selected. We compare two example selection strategies: (a) random selection and (b) diverse similarity-based selection. For both strategies, we observe a notable performance gain when $p$ and $q$ are set to 0.25 or higher, compared to settings where either $p$ or $q$ is 0 and all attributes are included. The gains are more pronounced in the random selection setting, where attribute filtering plays a greater role in guiding the model. In the first setting (Table 7(a)), where examples are selected randomly, we observe that using $p=0$ and $q=0$ yields an exact match accuracy of 0.846 and ROUGE-1 score (shown in parentheses) of 0.883. These scores represent the baseline performance without applying our method, and match the results of the FMW. As we increase $p$ and $q$ to 0.25, the performance improves significantly, reaching approximately 0.960 accuracy and 0.978 ROUGE-1. This indicates that even a modest level of filtering can effectively remove distracting features and highly improve model performance. As $p$ and $q$ increase further, performance remains consistently high. Notably, even at $p=0.75$ and $q=0.75$, where the model uses only 2 attributes, reducing nearly 95% of the input content in terms of characters (as discussed next), the accuracy remains close to 0.960. This shows that our method can retain the critical information required for imputation while drastically reducing the input size.
In the second setting (Table 7(b)), where tuples are selected using diverse similarity, the overall performance improves compared to random selection across all values of $p$ and $q$. The accuracy under $p=0$ and $q=0$ reaches 0.944, with a ROUGE-1 score of 0.951, showing that selecting diverse and informative examples enhances performance even without attribute filtering. As $p$ and $q$ increase, performance continues to improve. The best results are achieved at $p=0.75$, $q=0.50$ and $p=0.50$, $q=0.75$, where the model reaches an accuracy of 0.978 and ROUGE-1 scores of 0.990 and 0.986, respectively. However, these settings require more attributes compared to the setting $p=0.75$, $q=0.75$, which uses just 2 attributes (see Table 8) and achieves a very similar performance. Given that the difference in performance is minimal, the stricter setting with fewer attributes is a more practical choice, as it reduces input size while maintaining nearly identical accuracy. This can highlight the flexibility of our method in allowing users to balance performance with input efficiency.
To study the effect of varying p and q on attribute reduction, we observe that as the values of p and q increase, the algorithm becomes stricter in its selection. The results are shown in Table 8. When either parameter is set to 0, the method selects all 16 attributes. As both parameters increase, the average number of selected attributes steadily decreases. For instance, at $p=0.50$, $q=0.50$, only 7.6 attributes are selected on average, and this number drops to just 2 when $p=0.75$ and $q=0.75$. This trend confirms that tighter dependency thresholds are effective in filtering out weak or noisy attribute relationships, enabling the model to focus on the most informative signals.
Table 7. Effect of varying $(p,q)$ on imputation performance in the Zomato dataset with $k=3$
(a) Example selection: Random
[…]
0 0.846 (0.883)* 0.846 (0.883) 0.846 (0.883) 0.846 (0.883)
0.25 0.846 (0.883) 0.960 (0.978) 0.960 (0.977) 0.954 (0.974)
0.50 0.846 (0.883) 0.966 (0.982) 0.960 (0.973) 0.960 (0.975)
0.75 0.846 (0.883) 0.954 (0.972) 0.970 (0.985) 0.960 (0.980)
(b) Example selection: Diverse similarity
[…]
0 0.944 (0.951) 0.944 (0.951) 0.944 (0.951) 0.944 (0.951)
0.25 0.944 (0.951) 0.966 (0.979) 0.968 (0.978) 0.966 (0.977)
0.50 0.944 (0.951) 0.964 (0.976) 0.966 (0.978) 0.978 (0.990)
0.75 0.944 (0.951) 0.964 (0.976) 0.978 (0.986) 0.976 (0.983)
* Equivalent to the FMW result.
Equivalent to the LDI result.
Table 8. Effect of varying $(p,q)$ on the average # selected Attr. in the Zomato dataset
[…]
0 16 16 16 16
0.25 16 11.8 10.2 6.8
0.50 16 11.4 7.6 3.4
0.75 16 7.4 5 2
6.9. Limited-Resource Setting
To assess the effectiveness of LDI in resource-constrained environments, we evaluate its performance using Llama 3.2 3B, a smaller, open-source language model with significantly fewer parameters than GPT-4o-mini. This experiment compares LDI with FMW, which uses all attributes and randomly selected examples for imputation, without any filtering.
As shown in Table 9, LDI consistently outperforms the baseline across all datasets. These results show that LDI’s improvements are not dependent on large-scale language models; rather, its design effectively enhances performance through better input selection by providing the model with cleaner, more focused input. When smaller LLMs such as Llama 3.2 3B are used, focusing on localized input becomes especially important as these models have limited capacity to process and reason over lengthy inputs. If irrelevant attributes or uninformative examples are included, they can dilute the clues and distract the model from important patterns.
A particularly interesting case is the Zomato dataset with $k=10$, where there is a large drop in performance for both LDI and FMW. This dataset contains many lengthy attributes, and including 10 examples results in a large input prompt. In such cases, the model struggles to extract the patterns. Interestingly, using only 3 examples yields much better performance, suggesting that more examples do not always lead to better results when the context is noisy or overloaded. Even in this challenging setting, LDI performs considerably better than the baseline (0.252 vs. 0.128 in accuracy, and 0.404 vs. 0.286 in ROUGE F1), demonstrating its advantage in isolating informative content. These findings illustrate that guiding the model with concise, high-quality input, rather than simply more input, is key to achieving high performance.
Table 9. Performance comparison when deployed with a smaller language model (Llama 3.2 3B)
(a) $k=3$
Exact Match ROUGE-1 F1
Dataset FMW LDI Gain FMW LDI Gain
Buy 0.918 0.954 +3.9% 0.954 0.980 +2.7%
Restaurant 0.701 0.775 +10.6% 0.802 0.837 +4.4%
Zomato 0.766 0.894 +16.7% 0.826 0.928 +12.3%
Phone 0.910 0.966 +6.2% 0.941 0.977 +3.8%
(b) $k=10$
Exact Match ROUGE-1 F1
Dataset FMW LDI Gain FMW LDI Gain
Buy 0.923 0.944 +2.3% 0.959 0.973 +1.5%
Restaurant 0.738 0.770 +4.3% 0.833 0.839 +0.7%
Zomato 0.128 0.252 +96.9% 0.286 0.404 +41.3%
Phone 0.854 0.964 +12.9% 0.890 0.977 +9.8%
7. Limitations
While our approach effectively captures meaningful dependencies for imputation, it also has some limitations. First, the current implementation supports textual and categorical attributes and treats numerical values as textual tokens. This design choice was made because detecting patterns embedded in textual fields was the main challenge, requiring specialized handling of linguistic and symbolic variations. As a result, purely numerical relationships (e.g., arithmetic dependencies) are not explicitly modeled. Second, the method discovers dependencies through repetition patterns among tuples. This behavior is not unique to our framework, as all FD-based methods inherently depend on repeated evidence to infer dependencies. Nonetheless, we further introduced a group sampling strategy to relax this requirement, allowing the discovery of dependencies in cases where repetition is limited.
8. Conclusion and Future Work
In this paper, we proposed a localized imputation method that improves both the accuracy and explainability of LLM-based data imputation. Instead of relying on all available data or randomly selected samples, our approach selects a small set of informative attributes and representative tuples for each missing value. This focused design helps reduce noise, handle inconsistencies in real-world data, and guide the LLM toward more accurate predictions. Additionally, we make the imputation process more transparent by limiting the parts of the input that influence the prediction and explaining why they were chosen. Experiments on four real-world datasets show that our method achieves state-of-the-art performance while offering better explainability than existing LLM-based approaches.
This work opens several directions for future research: (1) One direction is to explore more advanced strategies for selecting informative attributes and tuples, particularly in cases where dependencies are conditional or involve more complex patterns. (2) Another direction is to explore transfer learning to adapt dependencies learned from one dataset to similar domains. (3) Future work can also enhance model explainability by going beyond limiting the input and identifying dependent parts, and instead, develop richer explanations that show how the model reasons about the data and arrives at each imputation.
Acknowledgements.
This research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). Additionally, Soroush Omidvartehrani is supported by the Alberta Innovates Graduate Student Scholarship (AI GSS).
Artifacts
The artifacts required to reproduce our experiments are publicly available at:
The repository provides the source code, datasets, and detailed instruction for experiment reproduction.